what is the answer to this problem? i have gotten stuck.

the square root of two thirds and the square root of three fourths.

sorry i forgot to say i needed it simplified.

studying for a chapter test tomorrow.

Answer by RAY100(1637)

sqrt (2/3),,,,,,,sqrt (3/4)
.
to simplify (reduce) we factor, but with perfect squares(1,4,9,16,25,36,49,etc)
.
Let's do sqrt (3/4) first,,,
.
This also equals ,,,,(sqrt 3)/ (sqrt 4),,,,
.
the sqrt of 3 has only 1 as a potential factor, and this does not help
The sqrt of 4 has 4 as a perfect square (2*2), and can be simplified
.
To (sqrt 3)/ (2),,, (just 2,, no sqrt),,,so
.
sqrt(3/4) = (sqrt 3) /2 = (1/2) sqrt 3
.
checking with a calc, sqrt(3/4) = .866,,,,1/2 sqrt3=.866,,,,ok
.
.
Returning to sqrt(2/3), there are no perfect square factors that will help, so this is in simplified form already.
.
sqrt(2/3) = sqrt .667 = .816
.
a Better example might be sqrt50 = sqrt(25*2)= sqrt25 * sqrt2= 5sqrt2
.
or sqrt 200 = sqrt(100*2) = sqrt(100) * sqrt2 = 10sqrt2
.